OpenStudy (anonymous):

solve using the elimination method 4x+3y=55 and 5x-3y=35

OpenStudy (anonymous):

4x+3y=55 5x-3y=35 --------- 9x =90 x =10 4(10)+3y=55 3y = 15 y = 3

OpenStudy (anonymous):

when i check it on the calculator x is correct but y is false. it says it should be 5 :S

OpenStudy (anonymous):

Sorry about that. 3y =15 y =5 not 3 as I put forth in the answer.